OpenStudy (anonymous):

The ordered pair (1, -5) is a solution for the system x + 2y = -9 and x - y = 6. True False

OpenStudy (swag):

false

OpenStudy (lgbasallote):

plug in the values and verify if the equations are correct.

OpenStudy (lgbasallote):

(1) + 2(-5) = -9 is it equal to -9? 1 - 10 = -9 -9 = -9 the first equation is correct...try the second one 1 - (-5) = 6 is it equal to -6? 1 + 5 = 6 6=6 the second equation is correct too...therefore the answer is TRUE